Solve by factoring: A rectangular painting with a width of x centimeters has an area of x^+50x square centimeters. Find a binomial that represents the length.
Answer by richwmiller(17219)

assuming you mean the area is x^2+50x
L*W=x^2+50x
L*x=x*(x+50)
L=x+50